OpenStudy (anonymous):

3y + 5 ≥ 2y + 1 > y - 1

OpenStudy (anonymous):

\[3y+5\geq 2y+1>y-1\] \[2y+5\geq y+1>-1\] so we have two inequalities to solve \[y+1>-1\] means \[y>-2\] and also \[2y+5\geq y+1\] \[y\geq -4\] putting these together tells you \[y>-2\]